# Competition of $\chi_{c}(2P)$ quarkonia and continuum in $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$

### v1

@@ IV Conclusions
In the present paper, we have concentrated on the possible observation of excited $P$-wave charmonia in the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction. One of our goals was to estimate the cross-section of meson-pair continua for both $D^{0}{\bar{D}}^{0}$ and $D^{+}D^{-}$ channels separately. We have considered $t$- and $u$-channel exchanges of the vector $D^{*,0}$ and $D^{*,\pm}$ mesons. The corresponding coupling constants for the $D^{*}D\gamma$ couplings have been obtained from the experimental radiative partial decay widths of $D^{*}\to D\gamma$. For the $D^{+}D^{-}$ channel, we have also considered the continuum mechanism via the $D^{\pm}$ $t/u$-channel exchanges plus contact term. Comparing our model results with the experimental data (Belle [4], BaBar [5]) we find a large contribution of the continuum to the $D^{0}{\bar{D}}^{0}$ channel and a much smaller one for the $D^{+}D^{-}$ channel (see Fig. 3).
Our analysis suggests that the broad bump at $M_{D\bar{D}}=3.8$ GeV observed in the $D^{0}{\bar{D}}^{0}$ channel, and not observed significantly in $D^{+}D^{-}$, is rather due to the continuum mechanism. However, a certain contribution of a broad resonance $\chi_{c0}(3860)$ cannot be completely ruled out. We have shown that taking this resonance into account with $\Gamma_{\chi_{c0}}\sim 50$ MeV improves the description of the near-threshold data.
@@
Acknowledgements.
The work of A.S. was partially supported by the Centre for Innovation and Transfer of Natural Sciences and Engineering Knowledge in Rzeszów (Poland).

### v2

@@ IV Conclusions
In the present paper, we have concentrated on the possible observation of excited $P$-wave charmonia in the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction. One of our goals was to estimate the cross-section of meson-pair continua for both $D^{0}{\bar{D}}^{0}$ and $D^{+}D^{-}$ channels separately. We have considered $t$- and $u$-channel exchanges of the vector $D^{*0}$ and $D^{*\pm}$ mesons. The corresponding coupling constants for the $D^{*}D\gamma$ couplings have been obtained from the experimental radiative partial decay widths of $D^{*}\to D\gamma$. For the $D^{+}D^{-}$ channel, we have also considered the continuum mechanism via the $D^{\pm}$ $t/u$-channel exchanges plus contact term. Comparing our model results with the experimental data (Belle [4], BaBar [5]) we find a large contribution of the continuum to the $D^{0}{\bar{D}}^{0}$ channel and a much smaller one for the $D^{+}D^{-}$ channel (see Fig. 3).
Our analysis suggests that the broad bump at $M_{D\bar{D}}=3.8$ GeV observed in the $D^{0}{\bar{D}}^{0}$ channel, and not observed significantly in $D^{+}D^{-}$, is rather due to the continuum mechanism. However, a certain contribution of a broad resonance $\chi_{c0}(3860)$ cannot be completely ruled out. We have shown that taking this resonance into account with $\Gamma_{\chi_{c0}}=50$ MeV improves the description of the near-threshold data, in particular, the BaBar data for the $D^{+}D^{-}$ channel shown in Fig. 3(b). It is rather difficult to understand a large difference between $\Gamma_{\chi_{c0}}$ and $\Gamma(\chi_{c0}\to D\bar{D})$ in the scenario with large $\Gamma_{\chi_{c0}}$ given in (2.24) which leads to surprisingly small branching fraction $B(\chi_{c0}\to D\bar{D})$. It should be noted that the potential-model predictions for the mass and width of the $\chi_{c0}(2P)$ state differ from the PDG averages. This indicates that further theoretical and experimental research is needed to understand properties of the $\chi_{c0}(2P)$ state more precisely.